OpenStudy (anonymous):

Preston read twenty-eight pages in 2 hours and Bethany read eighty-five pages in 5 hours. Which of the following statements is true?

OpenStudy (anonymous):

Can you post the statements?

OpenStudy (jlg030597):

Welcome to Open Study

OpenStudy (anonymous):

Preston:\[\frac{28pages}{2hours}=14\frac{pages}{hour}\] Bethany:\[\frac{85pages}{5hours}=17\frac{pages}{hour}\]

OpenStudy (anonymous):

Bethany can read more than Preston in 5 hours.

OpenStudy (texaschic101):

28/2 = 14 pages per hour (Preston) 85/5 = 17 pages per hour (Bethany)

OpenStudy (texaschic101):

Bethany reads faster